the length of a rectangle is 8 in more than twice its width. if the perimeter of the rectangle is 58 in, find the width of the rectangle
a)6 in
b)7 in
c)8 in
d)9 in

Answer by checkley75(3666)

L=2W+8
P=2L+2W
58=2(2W+8)+2W
58=4W+16+2W
58=6W+16
6W=58-16
6W=42
W=42/6
W=7 ANSWER IS C) 7 INCHES.
PROOF
58=2(2*7+8)+2*7
58=2(14+8)+14
58=2*22+14
58=44+14
58=58